OpenStudy (anonymous):

If tan x° =3/y and cos x° =y/z ,what is the value of tan x°?

OpenStudy (campbell_st):

can you check your question as you have posted a value for tan(x) and have been asked to find tan(x)

OpenStudy (anonymous):

Sorry. THe answers I was give to choose from are sin x=3/z sin x= 3y sin x= z/3 sin x=3z

OpenStudy (campbell_st):

ok... so I'll draw the triangle and label the sides 

OpenStudy (campbell_st):

I used the information you were given tan = opp/adj and cos = adj/hyp

OpenStudy (campbell_st):

you need to use sin = opp/hyp hope it helps

OpenStudy (anonymous):

Yes! Thank you!